# Latent Space Super-Resolution for Higher-Resolution Image Generation with Diffusion Models

### v1

@@ A.2 Analysis
Our qualitative results, presented in Figures J and K for 16$\times$ resolution and Figures L and M for 64$\times$ resolution, demonstrate consistent trends when comparing RGB upsampling variants to latent upsampling using our LSRNA framework. Specifically, RGB upsampling methods, whether based on interpolation or super-resolution, produce smoother images that lack fine details. In contrast, latent upsampling yields sharper and more detailed results.
Our quantitative results in Table D further confirm these observations that latent upsampling approaches (i.e., LSRNA-DemoFusion and LSRNA-Pixelsmith) consistently outperform their RGB upsampling counterparts. The improvements are particularly evident in patch-based metrics like pFID and pKID, which focus on capturing finer details. These results underscore the critical role of latent space upsampling in enhancing local detail fidelity and textures.
@@
Table D: RGB vs. Latent Space Upsampling on OpenImages-Valid ($\times$9). The best results marked in bold.
@@ C.3 Training
Table E: Quantitative comparison of image generation results by LSR training variants on OpenImages-Valid ($\times$9). All training is conducted on a single NVIDIA Tesla V100-SXM2 GPU, using SwinIR [33] and RCAN [62] as backbones with LIIF [8] as the upsampler. Based on a balance of training efficiency and performance, we adopt the v1 configuration.
@@
Quantitative results of image generation by various LSR training settings are presented in Table E, while also demonstrating the efficiency of the LSR training. Although the LSR is trained on the paired dataset constructed with relatively sparse downscaling factors, it can generalize to arbitrary scaling factors during inference, enabled by our multi-scale training scheme and the generalization capability of LIIF. Our motivation for using LIIF upsampler instead of traditional fixed-scale upsampler [51] lies in its ability to handle arbitrary resolutions with a single LSR module trained once.
Figure H: Qualitative results of RNA using Scharr edge map.
@@ D Edge Detection for RNA
To identify these areas, we consider using edge detection algorithms. However, we find that common edge detectors such as Scharr [49], LoG [35], and Gabor [14], which primarily focus on precise object boundaries, tend to produce artifacts such as overgeneration around contours or jagged contours when used as the basis for RNA. We present qualitative results in Figure H using the Scharr edge map, which is known for effectively capturing weak edges. As shown, while strong edges (e.g., on the train’s window) are sparsely detected, weak edges (e.g., on the tree) appear with excessively low intensity. This results in artifacts or over-enhanced details around strong edges when RNA is applied.
To address this, we adopt Canny edge detection [5], which allows us to prioritize weak edges by adjusting the lower and upper thresholds. By doing so, we can detect detailed regions rather than strictly connected edge lines, as demonstrated in Figure H. This region-based detection allows RNA to enhance local details effectively without introducing edge-based artifacts.
Table F: Quantitative comparison of image generation results by Canny edge detection thresholds on OpenImages-Valid (×$\times 9$) with DemoFusion. $e_{max}$ is set to 1.2 (default).
@@
Figure I: Ablation study of UNA (Uniform Noise Addition) strength on OpenImages-Valid (×$\times 9$) with DemoFusion. Dotted line shows our default RNA setting ($e_{min}=0$ and $e_{max}=1.2$).
Table G: Pixel-wise difference based on RNA strength. Differences are computed after applying RNA compared to no RNA. Histogram matching is applied before computing differences. The scene from the main Figure 7 is used.
@@ E Robustness of RNA
We demonstrate in Table F that generation performance is robust to variations in the Canny’s thresholds, as long as they are set to prioritize weak edges. In our implementation, we use a lower threshold of 0 and an upper threshold of 255. We further evaluate the robustness of RNA through additional experiments: Figure I indicates that RNA is quantitatively better than UNA (Uniform Noise Addition), while Table G shows that RNA indeed enhances details where the edge map is activated.
@@ F Additional Ablation Studies
Table H: LSR & RNA ablation on OpenImages-Valid ($\times$9). The best results marked in bold.
@@ F.1 Effectiveness of LSR & RNA
The results are summarized in Table H. For DemoFusion, incorporating the LSR module enhances performance by providing high-quality latent guidance, improving image generation quality even with fewer denoising steps and without progressive upscaling. The addition of the RNA module further boosts performance by enriching finer details and textures in the generated images.
@@ F.2 Impact of RNA Strength
Table I: Ablation Study on RNA strength with Pixelsmith on OpenImages-Valid ($\times$9). The best results marked in bold.
@@
Building on the RNA strength tuning results for LSRNA-DemoFusion presented in the main text, we further evaluate the impact of RNA strength on LSRNA-Pixelsmith, as shown in Table I. While LSRNA-DemoFusion achieves optimal performance with $e_{min}=0$ (and $e_{max}=1.2$), LSRNA-Pixelsmith performs best with $e_{min}=0.4$ and $e_{max}=0.8$. This difference in optimal RNA strength arises from the distinct roles played by the reference latent in the high-resolution generation process of each model. LSRNA-Pixelsmith likely requires a higher $e_{min}$ to ensure effective noise injection into the reference latent. The RNA strength determined from this validation process is consistently applied across all other experiments including those in the main text.
Figure J: RGB vs. Latent Space Upsampling for DemoFusion on 16$\times$. Prompt used is "the sun is setting over the ocean on a cloudy day". Best viewed ZOOMED-IN.
Figure K: RGB vs. Latent Space Upsampling for Pixelsmith on 16$\times$. Prompt used is "the sun is setting over the ocean on a cloudy day". Best viewed ZOOMED-IN.
Figure L: RGB vs. Latent Space Upsampling for DemoFusion on 64$\times$. Prompt used is "A mysterious forest with tall, ancient trees and beams of sunlight filtering through the mist, detailed moss-covered roots, 8k". Best viewed ZOOMED-IN.
Figure M: RGB vs. Latent Space Upsampling for Pixelsmith on 64$\times$. Prompt used is "A mysterious forest with tall, ancient trees and beams of sunlight filtering through the mist, detailed moss-covered roots, 8k". Best viewed ZOOMED-IN.

### v2

@@ A.2 Analysis
Our qualitative results, presented in Figures C and D for 16$\times$ resolution and Figures E and F for 64$\times$ resolution, demonstrate consistent trends when comparing RGB upsampling variants to latent upsampling using our LSRNA framework. Specifically, RGB upsampling methods, whether based on interpolation or super-resolution, produce smoother images that lack fine details. In contrast, latent upsampling yields sharper and more detailed results.
Our quantitative results in Table A further confirm these observations that latent upsampling approaches (i.e., LSRNA-DemoFusion and LSRNA-Pixelsmith) consistently outperform their RGB upsampling counterparts. The improvements are particularly evident in patch-based metrics like pFID and pKID, which focus on capturing finer details. These results underscore the critical role of latent space upsampling in enhancing local detail fidelity and textures.
@@
Table A: RGB vs. Latent Space Upsampling on OpenImages-Valid ($\times$9). The best results marked in bold.
@@ C.3 Training
Table B: Quantitative comparison of image generation results by LSR training variants on OpenImages-Valid ($\times$9). All training is conducted on a single NVIDIA Tesla V100-SXM2 GPU, using SwinIR [33] and RCAN [62] as backbones with LIIF [8] as the upsampler. Based on a balance of training efficiency and performance, we adopt the v1 configuration.
@@
Quantitative results of image generation by various LSR training settings are presented in Table B, while also demonstrating the efficiency of the LSR training. Although the LSR is trained on the paired dataset constructed with relatively sparse downscaling factors, it can generalize to arbitrary scaling factors during inference, enabled by our multi-scale training scheme and the generalization capability of LIIF. Our motivation for using LIIF upsampler instead of traditional fixed-scale upsampler [51] lies in its ability to handle arbitrary resolutions with a single LSR module trained once.
Figure A: Qualitative results of RNA using Scharr edge map.
@@ D Edge Detection for RNA
To identify these areas, we consider using edge detection algorithms. However, we find that common edge detectors such as Scharr [49], LoG [35], and Gabor [14], which primarily focus on precise object boundaries, tend to produce artifacts such as overgeneration around contours or jagged contours when used as the basis for RNA. We present qualitative results in Figure A using the Scharr edge map, which is known for effectively capturing weak edges. As shown, while strong edges (e.g., on the train’s window) are sparsely detected, weak edges (e.g., on the tree) appear with excessively low intensity. This results in artifacts or over-enhanced details around strong edges when RNA is applied.
To address this, we adopt Canny edge detection [5], which allows us to prioritize weak edges by adjusting the lower and upper thresholds. By doing so, we can detect detailed regions rather than strictly connected edge lines, as demonstrated in Figure A. This region-based detection allows RNA to enhance local details effectively without introducing edge-based artifacts.
Table C: Quantitative comparison of image generation results by Canny edge detection thresholds on OpenImages-Valid (×$\times 9$) with DemoFusion. $e_{max}$ is set to 1.2 (default).
@@
Figure B: Ablation study of UNA (Uniform Noise Addition) strength on OpenImages-Valid (×$\times 9$) with DemoFusion. Dotted line shows our default RNA setting ($e_{min}=0$ and $e_{max}=1.2$).
Table D: Pixel-wise difference based on RNA strength. Differences are computed after applying RNA compared to no RNA. Histogram matching is applied before computing differences. The scene from the main Figure 7 is used.
@@ E Robustness of RNA
We demonstrate in Table C that generation performance is robust to variations in the Canny’s thresholds, as long as they are set to prioritize weak edges. In our implementation, we use a lower threshold of 0 and an upper threshold of 255. We further evaluate the robustness of RNA through additional experiments: Figure B indicates that RNA is quantitatively better than UNA (Uniform Noise Addition), while Table D shows that RNA indeed enhances details where the edge map is activated.
@@ F Additional Ablation Studies
Table E: LSR & RNA ablation on OpenImages-Valid ($\times$9). The best results marked in bold.
@@ F.1 Effectiveness of LSR & RNA
The results are summarized in Table E. For DemoFusion, incorporating the LSR module enhances performance by providing high-quality latent guidance, improving image generation quality even with fewer denoising steps and without progressive upscaling. The addition of the RNA module further boosts performance by enriching finer details and textures in the generated images.
@@ F.2 Impact of RNA Strength
Table F: Ablation Study on RNA strength with Pixelsmith on OpenImages-Valid ($\times$9). The best results marked in bold.
@@
Building on the RNA strength tuning results for LSRNA-DemoFusion presented in the main text, we further evaluate the impact of RNA strength on LSRNA-Pixelsmith, as shown in Table F. While LSRNA-DemoFusion achieves optimal performance with $e_{min}=0$ (and $e_{max}=1.2$), LSRNA-Pixelsmith performs best with $e_{min}=0.4$ and $e_{max}=0.8$. This difference in optimal RNA strength arises from the distinct roles played by the reference latent in the high-resolution generation process of each model. LSRNA-Pixelsmith likely requires a higher $e_{min}$ to ensure effective noise injection into the reference latent. The RNA strength determined from this validation process is consistently applied across all other experiments including those in the main text.
Figure C: RGB vs. Latent Space Upsampling for DemoFusion on 16$\times$. Prompt used is "the sun is setting over the ocean on a cloudy day". Best viewed ZOOMED-IN.
Figure D: RGB vs. Latent Space Upsampling for Pixelsmith on 16$\times$. Prompt used is "the sun is setting over the ocean on a cloudy day". Best viewed ZOOMED-IN.
Figure E: RGB vs. Latent Space Upsampling for DemoFusion on 64$\times$. Prompt used is "A mysterious forest with tall, ancient trees and beams of sunlight filtering through the mist, detailed moss-covered roots, 8k". Best viewed ZOOMED-IN.
Figure F: RGB vs. Latent Space Upsampling for Pixelsmith on 64$\times$. Prompt used is "A mysterious forest with tall, ancient trees and beams of sunlight filtering through the mist, detailed moss-covered roots, 8k". Best viewed ZOOMED-IN.